OpenStudy (anonymous):

a line passes through points -10,-4 and -1,2. What is the y intercept of the line

OpenStudy (anonymous):

i got 8/3.

OpenStudy (anonymous):

its -16/9

OpenStudy (anonymous):

\[slope, m=\frac{ y-y1 }{ x-x1 }\] substituting the two points gives m=2/3 \[y=\frac{ 2 }{ 3 }x+c\] Substitute any of the points into x and y to find c=8/3.

OpenStudy (anonymous):

yeah! my bet